# LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation

### v1

@@ Abstract
Conditional layout generation aims to automatically generate visually appealing and semantically coherent layouts from user-defined constraints. While recent methods based on generative models have shown promising results, they typically require substantial amounts of training data or extensive fine-tuning, limiting their versatility and practical applicability. Alternatively, some training-free approaches leveraging in-context learning with Large Language Models (LLMs) have emerged, but they often suffer from limited reasoning capabilities and overly simplistic ranking mechanisms, which restrict their ability to generate consistently high-quality layouts. To this end, we propose LayoutCoT, a novel approach that leverages the reasoning capabilities of LLMs through a combination of Retrieval-Augmented Generation (RAG) and Chain-of-Thought (CoT) techniques. Specifically, LayoutCoT transforms layout representations into a standardized serialized format suitable for processing by LLMs. A Layout-aware RAG is used to facilitate effective retrieval and generate a coarse layout by LLMs. This preliminary layout, together with the selected exemplars, is then fed into a specially designed CoT reasoning module for iterative refinement, significantly enhancing both semantic coherence and visual quality. We conduct extensive experiments on five public datasets spanning three conditional layout generation tasks. Experimental results demonstrate that LayoutCoT achieves state-of-the-art performance without requiring training or fine-tuning. Notably, our CoT reasoning module enables standard LLMs, even those without explicit deep reasoning abilities, to outperform specialized deep-reasoning models such as deepseek-R1, highlighting the potential of our approach in unleashing the deep reasoning capabilities of LLMs for layout generation tasks.
footnotetext: Equal Contribution. Name is ordered by alphabet.
🖂Corresponding Author.
@@ 1 Introduction
Layout is a fundamental element in graphic design, determined by the deliberate and structured arrangement of visual components. Automated layout generation has recently emerged as an important research topic, aiming to significantly alleviate the workload of designers while addressing diverse user requirements [11, 19, 21, 22, 23, 26]. Current automated layout generation tasks broadly fall into three categories: i) content-aware layouts, which integrate visual or textual content for contextually appropriate designs (e.g., aligning text boxes with salient image regions in posters); ii) constraint-explicit layouts, which strictly adhere to predefined constraints, commonly found in user interface (UI) design and document typesetting; iii) and text-to-layout tasks, which translate abstract textual instructions into spatially coherent layouts, crucial for cross-modal creative tools.
@@
Early approaches [9, 12] mainly employed generative models, learning implicit design rules from extensive labeled datasets. Methods such as LayoutGAN [25] and LayoutDiffusion [43] have shown promising capabilities in generating plausible layouts, yet their reliance on substantial training data and domain-specific fine-tuning limits scalability and versatility. When faced with limited training data, purely generative methods [19, 16, 18, 14, 28] often fail to effectively capture sparse data distributions, restricting generalization and necessitating separate models for distinct tasks. Furthermore, the demand for extensive labeled datasets makes training generative layout models costly and inefficient, highlighting the necessity for a more flexible and versatile layout generation paradigm.
@@
With the advancement of large language models (LLMs) [2, 17, 1, 41, 42], layout generation methods are shifting towards a training-free paradigm. Recent research has shown that LLMs inherently possess a certain understanding of layout principles, such as alignment and spatial coherence [17, 3]. While leveraging this capability, approaches such as LayoutPrompter [28] utilize in-context learning combined with post-processing ranking mechanisms to guide LLMs in layout generation. However, this methodology demonstrates excessive focus on quantitative evaluation metrics, frequently resulting in structurally impractical bounding box configurations containing either disproportionately small elements or physically unfeasible arrangements. Moreover, when confronted with complex layout scenarios, the simplistic reasoning paradigms and elementary ranking systems inherent in such approaches tend to produce suboptimal compositions that violate both human visual perception principles and real-world application constraints.
@@
To unleash the deep reasoning abilities of LLMs in layout generation, we must resolve two key issues: first, how to select suitable contexts as references for layout generation; and second, how to design a chain of thought that enables LLMs to engage in human-like deep reasoning during layout design. Therefore, we propose LayoutCoT, a novel approach that integrates a layout-aware RAG with Chain-of-Thought (CoT) in LLMs to address those above limitations. Specifically, LayoutCoT employs a layout-aware exemplar retrieval mechanism, dynamically selecting suitable reference contexts using layout similarity [32], which better captures structural and semantic similarities compared to previous methods like LayoutPrompter [28]. We further introduce a meticulously designed Chain-of-Thought (CoT) reasoning framework that refines the layout through three sequential stages, each addressing a distinct aspect of the design process. In the first stage, the LLM is prompted to evaluate each element from the specified categories and determine their initial positions, emphasizing logical structure and visual appeal without accounting for interactions among element sizes. In the second stage, the LLM is guided to optimize the sizes and positions of these elements, effectively resolving overlaps or potential crowding issues. Finally, in the third stage, fine-grained mathematical adjustments are applied to specific layout dimensions to avoid unreasonable arrangements—such as the occlusion of salient elements—and to ensure both aesthetic quality and practical viability.
@@
We conducted comprehensive experiments across five public datasets (PKU [14], CGL [46], RICO [29], Publaynet [45], and WebUI [40]), spanning three distinct layout generation tasks. Our results demonstrate that LayoutCoT consistently outperforms existing state-of-the-art methods without requiring any training or fine-tuning. Notably, our CoT module enables standard LLMs, such as GPT-4, to surpass the performance of specialized deep-reasoning models like DeepSeek-R1 [10], underscoring the potential of our approach for practical layout generation tasks. The main contributions of our work are summarized as follows:
@@ 2 Related Work
Layout generation has emerged as an active area of research in recent years. It can be broadly divided into three categories: content-aware layout generation [14, 44], constraint-explicit layout generation [21, 22, 23], and text-to-layout generation [15, 27]. Content-aware layout generation aims to place layout elements on a given canvas in a visually pleasing manner while avoiding overlap with any background images on the canvas. Constraint-Explicit layout generation typically involves arranging predefined elements, or elements with specified dimensions, or even specified relationships among elements, on a blank interface in a reasonable configuration. Text-to-layout generation, meanwhile, focuses on creating layouts based on natural language descriptions, which is generally more challenging.
@@ 2.1 Generative layout generation
Early work in this field commonly utilized generative models for layout generation. Research based on Generative Adversarial Networks (GANs) [26, 14] and Variational Autoencoders (VAEs) [23, 20] pioneered the application of generative models to layout generation, achieving noticeable success in improving layout aesthetics through latent variable optimization. With the advent of the Transformer [38] architecture, numerous studies [11, 22, 15, 27, 33] emerged that adopted a sequence-to-sequence approach, further enhancing the quality of generated layouts. More recently, the proposal of diffusion [12] models has led to additional methods [16, 18] that bolster generative capabilities in scenarios with explicit layout constraints, offering further advancements in both quality and overall visual appeal. However, the existing generative layout methods are not capable of handling all types of layout generation tasks using a single model. Moreover, these approaches typically rely heavily on extensive datasets and require substantial training or fine-tuning, making practical applications challenging.
@@ 2.2 LLM-based Layout Generation
Large language models have demonstrated exceptional few-shot performance [2, 1, 6, 37] across various natural language processing (NLP) tasks. Recent studies have shown that LLMs themselves can generate layouts in structured descriptive languages such as HTML [28, 36]. For instance, LayoutNUWA [36] leverages Code Llama’s [34] code completion for layout generation, while PosterLlama [35] fine-tunes models such as MiniGPT-4 [48] to further enhance code completion capabilities for layout generation tasks. However, these methods rely heavily on sophisticated fine-tuning strategies and are dependent on large amounts of data. In contrast, LayoutPrompter [28] and LayoutGPT [7] utilize GPT-3 [2] via in-context learning to achieve training-free layout generation, revealing that LLMs inherently possess the capacity for layout generation. Although these methods do not require extensive, painstaking training and can generalize across various tasks, the quality of their generated layouts remains a concern.
@@ 2.3 Retrieval-Augmented Generation and Chain of Thought
Retrieval-Augmented Generation (RAG) [24] is a cutting-edge technique that combines retrieval and generation to enhance the model’s generative capabilities. By retrieving relevant examples or information [30, 5], RAG introduces more comprehensive contextual data during the generation process, thereby improving the quality of the outputs. Content-aware graphic layout generation aims to automatically arrange visual elements based on given content, such as e-commerce product images. In this context, the Retrieval-Augmented Layout Transformer (RALF) [13] leverages the retrieval of layout examples most similar to the input image to assist in layout design tasks. However, the RAG in RALF is constrained by its input paradigm, lacking plug-and-play characteristics and performing suboptimally across other tasks. Chain-of-Thought (CoT) [39, 4, 31] is a method that enhances model reasoning capabilities through the simulation of human thought processes. CoT achieves this by progressively solving problems through phased reasoning, offering more interpretable and coherent solutions to complex tasks. Although CoT has demonstrated powerful reasoning advantages across numerous domains, it has not yet been applied to layout design tasks. We attempt to integrate RAG with CoT for layout generation tasks, utilizing RAG to provide rich contextual information, while employing CoT for phased reasoning and optimization. This synergistic approach aims to enhance the quality and coherence of the generated results, combining the strengths of both methodologies effectively.
@@ 3.1 Layout-aware RAG
We employ a Retrieval-Augmented Generation (RAG) mechanism to dynamically select relevant layout examples as prompts, thereby enhancing the flexibility and coherence of generation. Specifically, let =$\mathcal{L}=\bigl{\{}(b_{i},c_{i})\bigr{\}}_{i=1}^{m}$ and ^=$\hat{\mathcal{L}}=\bigl{\{}(\hat{b}_{j},\hat{c}_{j})\bigr{\}}_{j=1}^{n}$ denote two layouts, where each element consists of a bounding box $b_{i}$ (or $\hat{b}_{j}$) and an associated label $c_{i}$ (or $\hat{c}_{j}$). We use a dissimilarity function from LTSim [32] $D(\mathcal{L},\hat{\mathcal{L}})$ based on a soft matching scheme:
$D(\mathcal{L},\hat{\mathcal{L}})\;=\;\min_{\Gamma}\sum_{(i,j)\in\Omega}\Gamma_{i,j}\,\mu\bigl{(}b_{i},\,c_{i},\;\hat{b}_{j},\,\hat{c}_{j}\bigr{)},\quad\text{s.t.}\;\Gamma_{i,j}\geq 0,$ (1)
@@
$\mathrm{LTSim}(\mathcal{L},\hat{\mathcal{L}})\;=\;\exp\Bigl{(}-\,D(\mathcal{L},\hat{\mathcal{L}})\Bigr{)}.$ (2)
A lower cost $D(\mathcal{L},\hat{\mathcal{L}})$ implies higher layout similarity; we set the scale parameter to $1.0$ for simplicity [32].
To retrieve the top-$K$ examples for a given layout $\mathcal{L}$, we consider a database $\mathbf{DB}=\{\hat{\mathcal{L}}_{1},\hat{\mathcal{L}}_{2},\dots,\hat{\mathcal{L}}_{N}\}$. We compute $\mathrm{LTSim}\bigl{(}\mathcal{L},\hat{\mathcal{L}}_{i}\bigr{)}$ for each candidate ^$\hat{\mathcal{L}}_{i}\in\mathbf{DB}$, and select the top-$K$ layouts that yield the highest scores:
$\mathcal{R}_{K}(\mathcal{L})\;=\;\mathrm{TopK}_{\,\hat{\mathcal{L}}\in\mathbf{DB}}\,\mathrm{LTSim}\bigl{(}\mathcal{L},\,\hat{\mathcal{L}}\bigr{)}.$ (3)
We then collect these ^$\hat{\mathcal{L}}_{k}\in\mathcal{R}_{K}(\mathcal{L})$ as prompting examples in subsequent generation stages. By leveraging the soft matching mechanism in $\mathrm{LTSim}$, the Layout-aware RAG remains robust even when the input and candidate layouts differ in the number or types of elements.

### v3

@@ Abstract
Conditional layout generation aims to automatically produce visually appealing and semantically coherent layouts under user-defined constraints. Generative models excel but require large task-specific datasets or costly fine-tuning, hampering scalability and real-world adoption. Training-free LLM methods, though cheaper, still suffer from limited reasoning depth and naive ranking strategies, leading to sub-optimal layouts. To this end, we introduce LayoutCoT, the first framework that explicitly unlocks the latent reasoning abilities of off-the-shelf LLMs for layout generation by uniting Layout-aware Retrieval-Augmented Generation (RAG) with Chain-of-Thought (CoT) reasoning. LayoutCoT first linearizes layouts into a LLM-friendly serialized form. Then it employs a layout-aware RAG module to retrieve relevant exemplars and then prompt the LLM to draft a coarse layout with the user’s constraints. Finally, we introduce a dedicated CoT module to iteratively refine this draft, reasoning step-by-step over spatial and semantic relations to produce layouts with markedly higher coherence and visual quality. On five public benchmarks covering three conditional layout generation tasks, LayoutCoT establishes new state-of-the-art results without fine-tuning. Remarkably, our CoT module enables a standard LLM to surpass specialized deep-reasoning models such as DeepSeek-R1, underscoring the promise of CoT-driven LLMs for scalable, training-free layout generation.
LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation
Hengyu Shithanks: Both authors contributed equally to this research. Names are ordered alphabetically., Junhao Su11footnotemark: 1, Tianyang Han, Junfeng Luo, Jialin Gaothanks: Corresponding author. Meituan {shihengyu02,sujunhao02,hantianyang,luojunfeng,gaojialin04}@meituan.com
@@ 1 Introduction
Layout is a fundamental element in graphic design, determined by the deliberate and structured arrangement of visual components. Automated layout generation has recently emerged as an important research topic, aiming to significantly alleviate the workload of designers while addressing diverse user requirements Gupta et al. (2021); Jiang et al. (2022); Kikuchi et al. (2021); Kong et al. (2022); Lee et al. (2020); Li et al. (2020). Current automated layout generation tasks broadly fall into three categories: i) content-aware layouts, which integrate visual or textual content for contextually appropriate designs (e.g., aligning text boxes with salient image regions in posters); ii) constraint-explicit layouts, which strictly adhere to predefined constraints, commonly found in user interface (UI) design and document typesetting; iii) and text-to-layout tasks, which translate abstract textual instructions into spatially coherent layouts, crucial for cross-modal creative tools.
@@
Early GAN and diffusion-based approaches Goodfellow et al. (2014); Ho et al. (2020) learned implicit design rules from extensive labeled datasets. Methods such as LayoutGAN Li et al. (2019) and LayoutDiffusion Zheng et al. (2023) have shown promising capabilities in generating plausible layouts, yet their reliance on substantial training data and domain-specific fine-tuning limits scalability and versatility. When faced with limited training data, purely generative methods Jiang et al. (2022); Hui et al. (2023); Inoue et al. (2023); Hsu et al. (2023); Lin et al. (2023b) often fail to effectively capture sparse data distributions, restricting generalization and necessitating separate models for distinct tasks. Furthermore, the demand for extensive labeled datasets makes training generative layout models costly and inefficient, highlighting the necessity for a more flexible and versatile layout generation paradigm.
@@
With the advancement of large language models (LLMs) Brown et al. (2020); Imani et al. (2024); Achiam et al. (2023); Yang et al. (2022a, b), layout generation methods are shifting towards a training-free paradigm. Recent research has shown that LLMs inherently possess a certain understanding of layout principles, such as alignment and spatial coherence Imani et al. (2024); Chen et al. (2021). While leveraging this capability, approaches such as LayoutPrompter Lin et al. (2023b) utilize in-context learning combined with post-processing ranking mechanisms to guide LLMs in layout generation. However, this methodology demonstrates excessive focus on quantitative evaluation metrics, frequently resulting in structurally impractical bounding box configurations containing either disproportionately small elements or physically unfeasible arrangements. Moreover, when confronted with complex layout scenarios, the simplistic reasoning paradigms and elementary ranking systems inherent in such approaches tend to produce suboptimal compositions that violate both human visual perception principles and real-world application constraints.
@@
To unleash the deep reasoning abilities of LLMs in layout generation, we must resolve two key issues: first, how to select suitable contexts as references for layout generation; and second, how to design a chain of thought that enables LLMs to engage in human-like deep reasoning during layout design. Therefore, we propose LayoutCoT, a novel approach that integrates a layout-aware RAG with Chain-of-Thought (CoT) in LLMs to address those above limitations. Specifically, LayoutCoT employs a layout-aware exemplar retrieval mechanism, dynamically selecting suitable reference contexts using layout similarity Otani et al. (2024), which better captures structural and semantic similarities compared to previous methods like LayoutPrompter Lin et al. (2023b). We further introduce a meticulously designed Chain-of-Thought (CoT) reasoning framework that refines the layout through three sequential stages, each addressing a distinct aspect of the design process. In the first stage, the LLM is prompted to evaluate each element from the specified categories and determine their initial positions, emphasizing logical structure and visual appeal without accounting for interactions among element sizes. In the second stage, the LLM is guided to optimize the sizes and positions of these elements, effectively resolving overlaps or potential crowding issues. Finally, in the third stage, fine-grained mathematical adjustments are applied to specific layout dimensions to avoid unreasonable arrangements—such as the occlusion of salient elements—and to ensure both aesthetic quality and practical viability.
@@
We conducted comprehensive experiments across five public datasets (PKU Hsu et al. (2023), CGL Zhou et al. (2022a), RICO Liu et al. (2018), Publaynet Zhong et al. (2019), and WebUI Wu et al. (2023)), spanning three distinct layout generation tasks. Our results demonstrate that LayoutCoT consistently outperforms existing state-of-the-art methods without requiring any training or fine-tuning. Notably, our CoT module enables standard LLMs, such as GPT-4, to surpass the performance of specialized deep-reasoning models like DeepSeek-R1 Guo et al. (2025), underscoring the potential of our approach for practical layout generation tasks. The main contributions of our work are summarized as follows:
@@ 2 Related Work
Layout generation has emerged as an active area of research in recent years. It can be broadly divided into three categories: content-aware layout generation Hsu et al. (2023); Zheng et al. (2019), constraint-explicit layout generation Kikuchi et al. (2021); Kong et al. (2022); Lee et al. (2020), and text-to-layout generation Huang et al. (2021); Lin et al. (2023a). Content-aware layout generation aims to place layout elements on a given canvas in a visually pleasing manner while avoiding overlap with any background images on the canvas. Constraint-Explicit layout generation typically involves arranging predefined elements, or elements with specified dimensions, or even specified relationships among elements, on a blank interface in a reasonable configuration. Text-to-layout generation, meanwhile, focuses on creating layouts based on natural language descriptions, which is generally more challenging.
@@ 2.1 Generative layout generation
Early work in this field commonly utilized generative models for layout generation. Research based on Generative Adversarial Networks (GANs) Li et al. (2020); Hsu et al. (2023) and Variational Autoencoders (VAEs) Lee et al. (2020); Jyothi et al. (2019) pioneered the application of generative models to layout generation, achieving noticeable success in improving layout aesthetics through latent variable optimization. With the advent of the Transformer Vaswani et al. (2017) architecture, numerous studies Gupta et al. (2021); Kong et al. (2022); Huang et al. (2021); Lin et al. (2023a); Rahman et al. (2021); Hu et al. (2025); Han et al. (2025) emerged that adopted a sequence-to-sequence approach, further enhancing the quality of generated layouts. More recently, the proposal of diffusion Ho et al. (2020) models has led to additional methods Hui et al. (2023); Inoue et al. (2023) that bolster generative capabilities in scenarios with explicit layout constraints, offering further advancements in both quality and overall visual appeal. However, the existing generative layout methods are not capable of handling all types of layout generation tasks using a single model. Moreover, these approaches typically rely heavily on extensive datasets and require substantial training or fine-tuning, making practical applications challenging.
@@ 2.2 LLM-based Layout Generation
Large language models have demonstrated exceptional few-shot performance Brown et al. (2020); Achiam et al. (2023); Chowdhery et al. (2023); Touvron et al. (2023) across various natural language processing (NLP) tasks. Recent studies have shown that LLMs themselves can generate layouts in structured descriptive languages such as HTML Lin et al. (2023b); Tang et al. (2023). For instance, LayoutNUWA Tang et al. (2023) leverages Code Llama’s Roziere et al. (2023) code completion for layout generation, while PosterLlama Seol et al. (2024) fine-tunes models such as MiniGPT-4 Zhu et al. (2023) to further enhance code completion capabilities for layout generation tasks. However, these methods rely heavily on sophisticated fine-tuning strategies and are dependent on large amounts of data. In contrast, LayoutPrompter Lin et al. (2023b) and LayoutGPT Feng et al. (2023) utilize GPT-3 Brown et al. (2020) via in-context learning to achieve training-free layout generation, revealing that LLMs inherently possess the capacity for layout generation. Although these methods do not require extensive, painstaking training and can generalize across various tasks, the quality of their generated layouts remains a concern.
@@ 2.3 Retrieval-Augmented Generation and Chain of Thought
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a cutting-edge technique that combines retrieval and generation to enhance the model’s generative capabilities. By retrieving relevant examples or information Liu et al. (2024); Cheng et al. (2025), RAG introduces more comprehensive contextual data during the generation process, thereby improving the quality of the outputs. Content-aware graphic layout generation aims to automatically arrange visual elements based on given content, such as e-commerce product images. In this context, the Retrieval-Augmented Layout Transformer (RALF) Horita et al. (2024) leverages the retrieval of layout examples most similar to the input image to assist in layout design tasks. However, the RAG in RALF is constrained by its input paradigm, lacking plug-and-play characteristics and performing suboptimally across other tasks. Chain-of-Thought (CoT) Wei et al. (2022); Chen et al. (2025); Luo et al. (2025) is a method that enhances model reasoning capabilities through the simulation of human thought processes. CoT achieves this by progressively solving problems through phased reasoning, offering more interpretable and coherent solutions to complex tasks. Although CoT has demonstrated powerful reasoning advantages across numerous domains, it has not yet been applied to layout design tasks. We attempt to integrate RAG with CoT for layout generation tasks, utilizing RAG to provide rich contextual information, while employing CoT for phased reasoning and optimization. This synergistic approach aims to enhance the quality and coherence of the generated results, combining the strengths of both methodologies effectively.
Table 1: Dataset statistics. Training Set refers to the candidate pool for RAG example selection.
Dataset Tasks Training Set Test Set Element Types
PKU content-aware 9,974 905 3
CGL content-aware 38,510 1,647 5
RICO constraint-explicit 31,694 3,729 25
PubLayNet constraint-explicit 311,397 10,998 5
WebUI text-to-layout 3,835 487 10
@@ 3 Methods
Figure 3: Details of the CoT Module. We illustrate the overall conversational logic of the multi-stage CoT. Depending on the type of task, there are slight variations in the details to adapt to specific requirements. Some of LayoutCoT prompts can be found in the supplementary material.
@@ 3.1 Layout-aware RAG
We employ a Retrieval-Augmented Generation (RAG) mechanism to dynamically select relevant layout examples as prompts, thereby enhancing the flexibility and coherence of generation. Specifically, let $\mathcal{L}=\bigl\{(b_{i},c_{i})\bigr\}_{i=1}^{m}$ and $\hat{\mathcal{L}}=\bigl\{(\hat{b}_{j},\hat{c}_{j})\bigr\}_{j=1}^{n}$ denote two layouts, where each element consists of a bounding box $b_{i}$ (or $\hat{b}_{j}$) and an associated label $c_{i}$ (or $\hat{c}_{j}$). We use a dissimilarity function from LTSim Otani et al. (2024) D(,^)D($\mathcal{L}$$\hat{\mathcal{L}}$) based on a soft matching scheme:
$D(\mathcal{L},\hat{\mathcal{L}})=\min_{\Gamma\geq 0}\sum_{(i,j)\in\Omega}\Gamma_{i,j}\,\mu\bigl(b_{i},c_{i},\hat{b}_{j},\hat{c}_{j}\bigr)$ (1)
@@
[…]
A lower cost D(,^)D($\mathcal{L}$$\hat{\mathcal{L}}$) implies higher layout similarity; we set the scale parameter to $1.0$ for simplicity Otani et al. (2024).
To retrieve the top-$K$ examples for a given layout $\mathcal{L}$, we consider a database $\mathbf{DB}=\{\hat{\mathcal{L}}_{1},\hat{\mathcal{L}}_{2},\dots,\hat{\mathcal{L}}_{N}\}$. We compute $\mathrm{LTSim}\bigl(\mathcal{L},\hat{\mathcal{L}}_{i}\bigr)$ for each candidate $\hat{\mathcal{L}}_{i}\in\mathbf{DB}$, and select the top-$K$ layouts that yield the highest scores:
[…]
We then collect these $\hat{\mathcal{L}}_{k}\in\mathcal{R}_{K}(\mathcal{L})$ as prompting examples in subsequent generation stages. By leveraging the soft matching mechanism in $\mathrm{LTSim}$, the Layout-aware RAG remains robust even when the input and candidate layouts differ in the number or types of elements.
Table 2: Quantitative comparison on Content-aware Layout Generation. We report results on both PKU (top) and CGL (bottom) datasets. LayoutCoT (Ours) consistently outperforms baselines across most metrics. Note that denotes variants using Gemini-2.5-pro, and denotes using DeepSeek-R1 without CoT.
Method Training-free Content Graphic
[…]
Dataset: PKU
CGL-GAN Zhou et al. (2022b) 0.219 0.175 0.226 0.0062 0.575 0.259 0.0605 0.707
DS-GAN Hsu et al. (2023) 0.209 0.217 0.229 0.0046 0.585 0.424 0.0260 0.903
RALF Horita et al. (2024) 0.171 0.150 0.246 0.0007 0.893 0.840 0.0180 0.998 0.981
LayoutPrompter Lin et al. (2023b) 0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997 0.865
LayoutPrompter Lin et al. (2023b) 0.184 0.179 0.215 0.0025 0.799 0.751 0.0011 0.995 0.869
LayoutCoT 0.214 0.180 0.264 0.0016 0.925 0.849 0.0004 1.000 0.806
LayoutCoT (ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000 1.000
LayoutCoT (ours) 0.143 0.139 0.218 0.0011 0.821 0.770 0.0024 1.000 1.000
Dataset: CGL
CGL-GAN Zhou et al. (2022b) 0.489 0.268 0.147 0.0420 0.275 0.244 0.269 0.876
DS-GAN Hsu et al. (2023) 0.451 0.224 0.193 0.0485 0.370 0.301 0.075 0.893
RALF Horita et al. (2024) 0.336 0.197 0.247 0.0023 0.943 0.884 0.027 0.995 0.874
LayoutPrompter Lin et al. (2023b) 0.415 0.184 0.227 0.0091 0.338 0.317 0.0049 0.996 0.812
LayoutCoT 0.267 0.182 0.225 0.0018 0.515 0.501 0.0023 0.994 0.803
LayoutCoT (ours) 0.170 0.101 0.260 0.0005 0.958 0.958 0.0018 0.998 0.939